OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 2, 6, 18, 54, ...

OpenStudy (anonymous):

A.)\[a _{n}= 3(2)^{n-1}\] B.)\[a _{n}=2(4)^{n-1}\] C.)\[a _{n}= 2(3)^{n-1}\]

OpenStudy (anonymous):

I think it's C but I'm not sure...

OpenStudy (psymon):

You're right.

OpenStudy (anonymous):

Cool, thanks :) Can you help me with another question on the same sequence?

OpenStudy (psymon):

What is it?

OpenStudy (anonymous):

Give the tenth term in the geometric sequence. 2, 6, 18, 54, ...

OpenStudy (psymon):

Now you're simply plugging in 10 for n in the equation you just found.

OpenStudy (anonymous):

Ohhh!!!! That's easy :p

OpenStudy (psymon):

Ikr? Not that bad, it's just unusual to see sequences in a formula form at first.

OpenStudy (anonymous):

Well thanks for your help :)

OpenStudy (psymon):

Sure